# Corona rigidity

### v3

@@ 7.5. Other models for rigidity
We continue the narrative of §7.1 by reviewing some of the consistency results regarding the rigidity questions for the Borel quotients of the Boolean algebra $\mathcal{P}({\mathbb{N}})$ and their generalizations to noncommutative $\mathrm{C}^{*}$-algebras. We will briefly discuss some iterated forcing models and, more specifically, the forcing extension used in [70], in which all the isomorphisms between the Borel quotients of $\mathcal{P}({\mathbb{N}})$ are topologically trivial. Then we review the main theorem of [78] which uses a slightly different forcing to generalize the result of [70] to the Borel quotients of direct products of sequences of matrix algebras. As we shall see, the results of [70] and [78] are more peculiar when noticing that in these generic extensions the Calkin algebra has outer automorphisms.
@@
The general idea is as in §7.1, with a few twists. The forcing extension used in [70] is a countable support iteration of length $\mathfrak{c}^{+}$ of creature forcings (see [134]) and random reals (see [7, §3.2]). The creature forcings assure that the ideal of sets on which $\Phi$ is topologically trivial is dense while the random reals facilitate the uniformization, namely that ${\mathbb{N}}$ is in this ideal. Several definability and absoluteness results from descriptive set theory have been used along the way in the proof. In particular, it is crucial that these forcings are proper, Suslin [^35] A forcing notion is Suslin if its underlying set is an analytic set of reals and both $\leq$ and $\bot$ are analytic relations. and ${\mathbb{N}}^{\mathbb{N}}$-bounding. The reflection argument used here is analogous to the one described at the beginning of §7.1.
The results of [70] were generalized in [78] to the noncommutative setting of the reduced products of the matrix algebras. Moreover, in the forcing extension used in [78] the creature forcings are replaced with a ‘groupwise Silver forcing’.
@@ 7.5.3. Reduced products of matrix algebras
This is a generalization of the algebras ${\mathcal{M}}_{X}$ introduced in Definition 5.21.
@@ 8. Endomorphisms
An example where the rigidity phenomena verified under Forcing Axioms for homeomorphisms extend to other classes of functions, arises in the context of continuous surjections between Čech–Stone remainders of locally compact, second countable, spaces. In [39], Dow and Hart proved that, if $\beta X\setminus X$ is a continuous image of $\beta{\mathbb{N}}\setminus{\mathbb{N}}$, then $X$ is the union of ${\mathbb{N}}$ with a compact set, when $\mathrm{OCA_{T}}$ is assumed. Their result was later extended in [170, Theorem E] as follows.
@@
[50, Theorem 4.9.1] Assume $\mathrm{OCA_{T}}$ and $\mathrm{MA}_{\aleph_{1}}$, and let $X$ and $Y$ be locally compact, second countable, zero-dimensional spaces. If in addition $X$ is countable, then for every continuous $f\colon\beta X\setminus X\to\beta Y\setminus Y$ there are a clopen partition $U\sqcup V=\beta X$ and a continuous $g\colon V\to\beta Y$ such that $f[U]$ is nowhere dense in $\beta Y\setminus Y$ and $f=g\restriction(\beta X\setminus X)$
The conclusion of Theorem 8.2 was called weak Extension Principle in [50].
Question 8.3.
Do $\mathrm{OCA_{T}}$ and $\mathrm{MA}$ imply the weak Extension Principle for every pair of locally compact, second countable spaces $X,Y$ and every continuous $f\colon\beta X\setminus X\to\beta Y\setminus Y$?
The conclusion of Theorem 8.2 looks much better with $U=\emptyset$, and in [50, §4.11] it was all but conjectured that this “Strong Extension Principle” follows from forcing axioms. Alas, the following example shows that we cannot take $U=\emptyset$.
@@
The techniques required to prove Theorem 8.6 are refinements of the tools developed to prove Theorem 7.12. Indeed a significant portion of the most technical results in [58] already applies to endomorphisms (the same arguments are also exposed in full detail in [59, §17]). By the appropriate analog of Lemma 7.14, it suffices to prove that the restriction of the endomorphism under consideration to some corner $p\mathcal{Q}(H)p$ is algebraically trivial.
Along with some elementary observations on endomorphisms of $\mathcal{B}(H)$ (see [164, Remark 2.3]), Theorem 8.6 permits to completely classify unital endomorphisms of $\mathcal{Q}(H)$ up to unitary equivalence via the Fredholm index of the image of the unilateral shift $s$. What can be obtained under $\mathrm{OCA_{T}}$ is thus a result in the same spirit of the early studies on essential unitary equivalence discussed in §4.
Theorem 8.7 ([164, Theorem 1.1]).
@@
A further difference in the behavior of the index map $\Theta$, when assuming CH, is its range. Under $\mathrm{OCA_{T}}$ the range of $\Theta$ coincides with the set of positive integers. On the other hand, $\mathrm{CH}$ entails the existence of a unital endomorphism $\Phi\in\operatorname{End}(\mathcal{Q}(H))$ that maps every unitary of $\mathcal{Q}(H)$ into a unitary with index zero, thus giving $\Theta([\Phi])=0$ (this follows from the main result of [66]; see also [164, Example 5.4]). It is not known whether the map $\Theta$ can assume negative values consistently with $\mathrm{ZFC}$. This problem relates to Question 4.16, since a $\Phi\in\operatorname{End}(\mathcal{Q}(H))$ such that $\Theta([\Phi])=-1$ is an endomorphism which, up to unitary equivalence, sends the unilateral shift to its adjoint. Finding such an endomorphism does not seem to be a much easier task than answering Question 4.16 itself.
@@ 9.0.1. Large Calkin algebras
in particular $\mathcal{K}(H)=\mathcal{K}_{\aleph_{0}}(H)$. It is not difficult to see that every norm-closed, self-adjoint ideal of $\mathcal{B}(H)$ is of this form for some $\lambda$ ([59, Proposition 12.3.4]). In particular, for any $\aleph_{\alpha}$ the proper ideals of $\mathcal{Q}(\ell_{2}(\aleph_{\alpha}))$ are linearly ordered by inclusion in type $\alpha$. Suppose that $\Phi$ is an automorphism of the Calkin algebra $\mathcal{Q}(H)=\mathcal{B}(H)/\mathcal{K}(H)$. Then $\Phi$ sends $\mathcal{K}_{\lambda}(H)/\mathcal{K}(H)$ to itself for every uncountable $\lambda\leq\kappa$. A moment of thought reveals that it is not at all clear whether the existence of an outer automorphism of $\mathcal{Q}(\ell_{2}(\aleph_{0}))$ implies the existence of an outer automorphism of $\mathcal{Q}(\ell_{2}(\kappa))$ for any uncountable value of $\kappa$. Indeed, it is an open problem whether the existence of an outer automorphism of $\mathcal{Q}(\ell_{2}(\kappa))$ is relatively consistent with $\mathrm{ZFC}$ for a single uncountable $\kappa$. Constructions of awkward objects using $\mathrm{CH}$ are typically easier (and historically obtained much earlier) than the proofs from forcing axioms that those objects do not exist. Nevertheless, a modification of Veličković’s argument can be adapted, by introducing curious objects called ‘Polish Aronszajn trees’, to show that $\mathrm{PFA}$ implies all automorphisms of $\mathcal{Q}(\ell_{2}(\kappa))$ are inner for all infinite $\kappa$ ([57]), in spite of an observation that the noncommutative analog of the poset for adding uncountably many Cohen reals fails to be ccc.
@@
The answer is easily positive if $\kappa=\lambda$ and $2^{\kappa}=\kappa^{+}$ ([68, Theorem 1]), but nothing else is known.
@@ 9.0.3. $\mathcal{P}(\kappa)/\operatorname{Fin}$ again
The study of automorphisms of $\mathcal{P}(\kappa)/\operatorname{Fin}$ for uncountable $\kappa$ is only loosely related to our story, but it hides surprises worth mentioning. If $\kappa$ is uncountable, then $\mathrm{MA}_{\aleph_{1}}$ and $\mathrm{OCA_{T}}$ already imply that all automorphisms of $\mathcal{P}(\kappa)/\operatorname{Fin}$ are trivial, for all $\kappa\leq\mathfrak{c}$ ([106, Theorem 3.4]). In [151, Corollary 1.2] this conclusion was extended to all $\kappa$ smaller than the first inaccessible cardinal (or all $\kappa$, if there are no inaccessible cardinals).
@@ 10. Uniform Roe coronas
To get a $\Phi_{*}$ as above, we need Forcing Axioms. Using techniques similar to the ones of §7.3, we show that under $\mathrm{OCA_{T}}$ and $\mathrm{MA}_{\aleph_{1}}$, $\Phi\restriction\ell_{\infty}(X)/c_{0}(X)$ must admit a Borel lifting $\ell_{\infty}(X)\to\mathrm{C}^{*}_{u}(Y)$. Consequently, using Ulam-stability type of results (in particular [111, Proposition 7.6 and 7.8]), we can show that the -homomorphism $\Phi\restriction\ell_{\infty}(X)/c_{0}(X)\to\mathrm{Q}^{*}_{u}(Y)$ is algebraically trivial (in the terminology of §8), and therefore there exists a -homomorphism $\ell_{\infty}(X)\to\mathrm{C}^{*}_{u}(Y)$ that lifts $\Phi$.
@@ 10.1. The Higson corona
Another interesting quotient $\mathrm{C}^{*}$-algebra associated to a u.l.f. metric space $(X,d)$ is the Higson corona. This algebra was introduced by Higson (yet it is sometimes referred to as the Higson–Roe corona) in connection with a $K$-theoretic analysis of the Roe index theorem for non-compact Riemannian manifolds. It is studied for its relevance in the connections between topology and coarse geometry ([42]) and index theory (e.g. [129, §6 and 7] and [174]). In general, the Higson corona gives a unique access to study the ends (or directions) of a u.l.f. metric space, and specifically of a finitely generated group.
@@
It is a good exercise (see [8, Proposition 3.6]) to show that $C_{\nu}(X)$ is exactly the center of the uniform Roe corona $\mathrm{Q}^{*}_{u}(X)$, hence two coarsely equivalent u.l.f. metric spaces have isomorphic Higson coronas. (This fact was proved without explicitly using the uniform Roe corona, see [130, §2.3]).
@@
Whoever was so far careful in going through this manuscript can now show, with some work of course, that under Forcing Axioms if $X$ is a u.l.f. metric space whose Higson corona is isomorphic to $\ell_{\infty}/c_{0}$, then $X$ must be coarsely equivalent to $\{n^{2}\mid n\in\omega\}$. A complete description of the situation in presence of Forcing Axioms was announced by the fourth author, who proved that, under $\mathrm{OCA_{T}}$ and $\mathrm{MA}_{\aleph_{1}}$, if two u.l.f. metric spaces $X$ and $Y$ have isomorphic Higson coronas, then $X$ and $Y$ are coarsely equivalent. This result has still to appear.
@@ 11.0.1. Dimension phenomena
For $n\geq 2$, the tensor product $\bigotimes_{n}\mathcal{P}({\mathbb{N}})$ can be identified with a Borel subalgebra of $\mathcal{P}({\mathbb{N}}^{n})$. The Stone dual of this tensor product is $(\beta{\mathbb{N}})^{n}$. Via this observation, the isomorphisms between finite powers of $\beta{\mathbb{N}}\setminus{\mathbb{N}}$ belong to our rigidity program. In [37], van Douwen proved that $(\beta{\mathbb{N}}\setminus{\mathbb{N}})^{m}$ and $(\beta{\mathbb{N}}\setminus{\mathbb{N}})^{n}$ are isomorphic if and only if $m=n$. This was extended in [53]. The most appealing formulation of this result is topological: for every continuous function $f$ from a product of compact Hausdorff spaces $\prod_{\xi<\kappa}X_{\xi}$ into $\beta{\mathbb{N}}$ [^43] The theorem is more general: it holds when $\beta{\mathbb{N}}$ is replaced by any $\beta{\mathbb{N}}$-space, as defined in [37]. there is a partition of the domain into clopen sets such that the restriction of $f$ to each one of them depends on at most one coordinate. This is one of the few nontrivial facts about $\beta{\mathbb{N}}\setminus{\mathbb{N}}$ that can be proved in $\mathrm{ZFC}$ (a similar result was proved from $\mathrm{OCA_{T}}$ in [91]). We do not know whether similar (dualised) statements hold for coronas of separable $\mathrm{C}^{*}$-algebras.
11.0.2. The role of MA and $\mathrm{OCA^{\#}}$

### v4

@@ 7.5. Other models for rigidity
We continue the narrative of §7.1 by reviewing some of the consistency results regarding the rigidity questions for the Borel quotients of the Boolean algebra $\mathcal{P}({\mathbb{N}})$ and their generalisations to noncommutative $\mathrm{C}^{*}$-algebras. We will briefly discuss some iterated forcing models and, more specifically, the forcing extension used in [74], in which all the isomorphisms between the Borel quotients of $\mathcal{P}({\mathbb{N}})$ are topologically trivial. Then we review the main theorem of [82] which uses a slightly different forcing to generalise the result of [74] to the Borel quotients of direct products of sequences of matrix algebras. As we shall see, the results of [74] and [82] are more peculiar when noticing that in these generic extensions the Calkin algebra has outer automorphisms.
@@
The general idea is as in §7.1, with a few twists. The forcing extension used in [74] is a countable support iteration of length $\mathfrak{c}^{+}$ of creature forcings (see [140]) and random reals (see [7, §3.2]). The creature forcings assure that the ideal of sets on which $\Phi$ is topologically trivial is dense while the random reals facilitate the uniformisation, namely that ${\mathbb{N}}$ is in this ideal. Several definability and absoluteness results from descriptive set theory have been used along the way in the proof. In particular, it is crucial that these forcings are proper, Suslin [^35] A forcing notion is Suslin if its underlying set is an analytic set of reals and both $\leq$ and $\bot$ are analytic relations. and ${\mathbb{N}}^{\mathbb{N}}$-bounding. The reflection argument used here is analogous to the one described at the beginning of §7.1.
The results of [74] were generalised in [82] to the noncommutative setting of the reduced products of the matrix algebras. Moreover, in the forcing extension used in [82] the creature forcings are replaced with a ‘groupwise Silver forcing’.
@@ 7.5.3. Reduced products of matrix algebras
This is a generalisation of the algebras ${\mathcal{M}}_{X}$ introduced in Definition 5.21.
@@
7.5.4. Large continuum
All models of $\mathrm{PFA}$ and all known models of $\mathrm{OCA_{T}}$ satisfy $\mathfrak{c}=\aleph_{2}$, and so do all models described in earlier sections. Models of $\mathrm{ZFC}$ with arbitrarily large continuum in which all automorphisms of $\mathcal{P}({\mathbb{N}})/\operatorname{Fin}$ are trivial were obtained in [97, p. 327] (using a modification of oracle-cc) and then again in [40], this time together with the $\mathrm{MA}$.
@@ 8. Endomorphisms
An example where the rigidity phenomena verified under Forcing Axioms for homeomorphisms extend to other classes of functions, arises in the context of continuous surjections between Čech–Stone remainders of locally compact, second countable, spaces. In [41], Dow and Hart proved that, if $\beta X\setminus X$ is a continuous image of $\beta{\mathbb{N}}\setminus{\mathbb{N}}$, then $X$ is the union of ${\mathbb{N}}$ with a compact set, when $\mathrm{OCA_{T}}$ is assumed. Their result was later extended in [178, Theorem E] as follows.
@@
[52, Theorem 4.9.1] Assume $\mathrm{OCA_{T}}$ and $\mathrm{MA}_{\aleph_{1}}$, and let $X$ and $Y$ be locally compact, second countable, zero-dimensional spaces. If in addition $X$ is countable, then for every continuous $f\colon\beta X\setminus X\to\beta Y\setminus Y$ there are a clopen partition $U\sqcup V=\beta X$ and a continuous $g\colon V\to\beta Y$ such that $f[U]$ is nowhere dense in $\beta Y\setminus Y$ and $f=g\restriction(\beta X\setminus X)$
This theorem was stated and proved for arbitrary finite powers of $\beta X\setminus X$ and $\beta Y\setminus Y$. Soon afterwards it was discovered that the above implies the full version of the theorem because continuous maps between Čech–Stone remainders of locally compact second countable spaces respect coordinates (see §11.0.1). By these results, Definition 8.3 and Theorem 8.4 below are equivalent to their analogues stated for arbitrary finite powers.
The conclusion of Theorem 8.2 was called weak Extension Principle in [52]. This principle was extended beyond the zero-dimensional setting in [179]. Its (slightly) more intricate definition is due to the fact that clopen sets in $\beta X\setminus X$ do not necessarily come from clopen sets in $\beta X$ when $X$ is not zero-dimensional.
Definition 8.3.
Let $X$ and $Y$ be locally compact noncompact second countable topological spaces. We say that $X$ and $Y$ satisfy the weak Extension Principle if the following happens:
For every continuous map $F\colon\beta X\setminus X\to\beta Y\setminus Y$ there exists a partition into clopen sets $\beta X\setminus X=U\cup V$, an open set with compact closure $V_{X}\subseteq X$ such that $F[U]$ is nowhere dense in $\beta Y\setminus Y$, and a continuous function $G\colon\beta(X\setminus V_{X})\to\beta Y$ which restricts to $F$ on $V$.
The weak Extension Principle is the statement “All pairs of locally compact noncompact second countable topological spaces satisfy the weak Extension Principle”.
The following is the main result of [179].
Theorem 8.4.
Assume $\mathrm{OCA_{T}}$ and $\mathrm{MA}_{\aleph_{1}}$. Then the weak Extension Principle holds.
The conclusions of Theorem 8.2 and Definition 8.3 look much better with $U=\emptyset$, and in [52, §4.11] it was all but conjectured that this “Strong Extension Principle” follows from forcing axioms. Alas, the following example shows that we cannot take $U=\emptyset$.
@@
The techniques required to prove Theorem 8.7 are refinements of the tools developed to prove Theorem 7.12. Indeed a significant portion of the most technical results in [60] already applies to endomorphisms (the same arguments are also exposed in full detail in [61, §17]). By the appropriate analog of Lemma 7.14, it suffices to prove that the restriction of the endomorphism under consideration to some corner $p\mathcal{Q}(H)p$ is algebraically trivial.
Along with some elementary observations on endomorphisms of $\mathcal{B}(H)$ (see [171, Remark 2.3]), Theorem 8.7 permits to completely classify unital endomorphisms of $\mathcal{Q}(H)$ up to unitary equivalence via the Fredholm index of the image of the unilateral shift $s$. What can be obtained under $\mathrm{OCA_{T}}$ is thus a result in the same spirit of the early studies on essential unitary equivalence discussed in §4.
Theorem 8.8 ([171, Theorem 1.1]).
@@
A further difference in the behavior of the index map $\Theta$, when assuming CH, is its range. Under $\mathrm{OCA_{T}}$ the range of $\Theta$ coincides with the set of positive integers. On the other hand, $\mathrm{CH}$ entails the existence of a unital endomorphism $\Phi\in\operatorname{End}(\mathcal{Q}(H))$ that maps every unitary of $\mathcal{Q}(H)$ into a unitary with index zero, thus giving $\Theta([\Phi])=0$ (this follows from the main result of [70]; see also [171, Example 5.4]). It is not known whether the map $\Theta$ can assume negative values consistently with $\mathrm{ZFC}$. This problem relates to Question 4.16, since a $\Phi\in\operatorname{End}(\mathcal{Q}(H))$ such that $\Theta([\Phi])=-1$ is an endomorphism which, up to unitary equivalence, sends the unilateral shift to its adjoint. Finding such an endomorphism does not seem to be a much easier task than answering Question 4.16 itself.
@@ 9.0.1. Large Calkin algebras
in particular $\mathcal{K}(H)=\mathcal{K}_{\aleph_{0}}(H)$. It is not difficult to see that every norm-closed, self-adjoint ideal of $\mathcal{B}(H)$ is of this form for some $\lambda$ ([61, Proposition 12.3.4]). In particular, for any $\aleph_{\alpha}$ the proper ideals of $\mathcal{Q}(\ell_{2}(\aleph_{\alpha}))$ are linearly ordered by inclusion in type $\alpha$. Suppose that $\Phi$ is an automorphism of the Calkin algebra $\mathcal{Q}(H)=\mathcal{B}(H)/\mathcal{K}(H)$. Then $\Phi$ sends $\mathcal{K}_{\lambda}(H)/\mathcal{K}(H)$ to itself for every uncountable $\lambda\leq\kappa$. A moment of thought reveals that it is not at all clear whether the existence of an outer automorphism of $\mathcal{Q}(\ell_{2}(\aleph_{0}))$ implies the existence of an outer automorphism of $\mathcal{Q}(\ell_{2}(\kappa))$ for any uncountable value of $\kappa$. Indeed, it is an open problem whether the existence of an outer automorphism of $\mathcal{Q}(\ell_{2}(\kappa))$ is relatively consistent with $\mathrm{ZFC}$ for a single uncountable $\kappa$. Constructions of awkward objects using $\mathrm{CH}$ are typically easier (and historically obtained much earlier) than the proofs from forcing axioms that those objects do not exist. Nevertheless, a modification of Veličković’s argument can be adapted, by introducing curious objects called ‘Polish Aronszajn trees’, to show that $\mathrm{PFA}$ implies all automorphisms of $\mathcal{Q}(\ell_{2}(\kappa))$ are inner for all infinite $\kappa$ ([59]), in spite of an observation that the noncommutative analog of the poset for adding uncountably many Cohen reals fails to be ccc.
@@
The answer is easily positive if $\kappa=\lambda$ and $2^{\kappa}=\kappa^{+}$ ([72, Theorem 1]), but nothing else is known.
@@ 9.0.3. $\mathcal{P}(\kappa)/\operatorname{Fin}$ again
The study of automorphisms of $\mathcal{P}(\kappa)/\operatorname{Fin}$ for uncountable $\kappa$ is only loosely related to our story, but it hides surprises worth mentioning. If $\kappa$ is uncountable, then $\mathrm{MA}_{\aleph_{1}}$ and $\mathrm{OCA_{T}}$ already imply that all automorphisms of $\mathcal{P}(\kappa)/\operatorname{Fin}$ are trivial, for all $\kappa\leq\mathfrak{c}$ ([112, Theorem 3.4]). In [157, Corollary 1.2] this conclusion was extended to all $\kappa$ smaller than the first inaccessible cardinal (or all $\kappa$, if there are no inaccessible cardinals).
Rigidity for the quotient of $\mathcal{P}(\kappa)$ modulo the ideal of subsets of cardinality smaller than $\kappa$ was studied in [112, 108, 107].
@@ 10. Uniform Roe coronas
To get a $\Phi_{*}$ as above, we need Forcing Axioms. Using techniques similar to the ones of §7.3, we show that under $\mathrm{OCA_{T}}$ and $\mathrm{MA}_{\aleph_{1}}$, $\Phi\restriction\ell_{\infty}(X)/c_{0}(X)$ must admit a Borel lifting $\ell_{\infty}(X)\to\mathrm{C}^{*}_{u}(Y)$. Consequently, using Ulam-stability type of results (in particular [117, Proposition 7.6 and 7.8]), we can show that the -homomorphism $\Phi\restriction\ell_{\infty}(X)/c_{0}(X)\to\mathrm{Q}^{*}_{u}(Y)$ is algebraically trivial (in the terminology of §8), and therefore there exists a -homomorphism $\ell_{\infty}(X)\to\mathrm{C}^{*}_{u}(Y)$ that lifts $\Phi$.
@@ 10.1. The Higson corona
Another interesting quotient $\mathrm{C}^{*}$-algebra associated to a u.l.f. metric space $(X,d)$ is the Higson corona. This algebra was introduced by Higson (yet it is sometimes referred to as the Higson–Roe corona) in connection with a $K$-theoretic analysis of the Roe index theorem for non-compact Riemannian manifolds. It is studied for its relevance in the connections between topology and coarse geometry ([44]) and index theory (e.g. [135, §6 and 7] and [183]). In general, the Higson corona gives a unique access to study the ends (or directions) of a u.l.f. metric space, and specifically of a finitely generated group.
@@
The Higson remainder $\nu X$ is the spectrum of the Higson corona, that is, the compact topological space such that $C_{\nu}(X)\cong C(\nu X)$. It is a good exercise (see [8, Proposition 3.6]) to show that $C_{\nu}(X)$ is exactly the center of the uniform Roe corona $\mathrm{Q}^{*}_{u}(X)$.
The following is (a restricted version of) Proposition 2.41 in [136].
Proposition 10.5.
Let $X$ and $Y$ be infinite u.l.f. metric spaces. Every proper [^42] In the setting of u.l.f. spaces properness reduces to ‘finite-to-one’. coarse map $\varphi\colon X\to Y$ extends to a continuous map $\nu\varphi\colon\nu X\to\nu Y$. Also, $\varphi$ is a coarse equivalence if and only if $\nu\varphi$ is a homeomorphism.
@@
What about the other side of the coin? The problem of establishing rigidity of Higson coronas under suitable Forcing Axioms was studied by the fourth author, who in [176] proved the following:
Theorem 10.7.
Assume $\mathrm{OCA_{T}}$ and $\mathrm{MA}_{\aleph_{1}}$. Let $X$ and $Y$ be u.l.f. metric spaces. If $X$ and $Y$ have isomorphic Higson coronas then they are coarsely equivalent.
In addition to Theorem 10.7, [176] characterises in a strong way all -homomorphisms between Higson coronas, stating the coarse version of the weak Extension Principle (see §8) and proving it in presence of Forcing Axioms.
@@ 11.0.1. Dimension phenomena
For $n\geq 2$, the tensor product $\bigotimes_{n}\mathcal{P}({\mathbb{N}})$ can be identified with a Borel subalgebra of $\mathcal{P}({\mathbb{N}}^{n})$. The Stone dual of this tensor product is $(\beta{\mathbb{N}})^{n}$. Via this observation, the isomorphisms between finite powers of $\beta{\mathbb{N}}\setminus{\mathbb{N}}$ dualise to isomorphisms between quotients ${\mathcal{M}}/{\mathcal{I}}$ where ${\mathcal{M}}$ is a Borel subalgebra of $\mathcal{P}({\mathbb{N}})$ and ${\mathcal{I}}$ is a Borel ideal, and hence belong to our rigidity program. In [38], van Douwen proved that $(\beta{\mathbb{N}}\setminus{\mathbb{N}})^{m}$ and $(\beta{\mathbb{N}}\setminus{\mathbb{N}})^{n}$ are isomorphic if and only if $m=n$. This was extended in [55]. The most appealing formulation of this result is topological: for every continuous function $f$ from a product of compact Hausdorff spaces $\prod_{\xi<\kappa}X_{\xi}$ into $\beta{\mathbb{N}}$ [^44] The theorem is more general: it holds when $\beta{\mathbb{N}}$ is replaced by any $\beta{\mathbb{N}}$-space, as defined in [38]. there is a partition of the domain into clopen sets such that the restriction of $f$ to each one of them depends on at most one coordinate (a similar result was proved from $\mathrm{OCA_{T}}$ in [95]). This is one of the few nontrivial facts about $\beta{\mathbb{N}}\setminus{\mathbb{N}}$ that can be proved in $\mathrm{ZFC}$. The only known result along these lines for $\mathrm{C}^{*}$-algebras is that no corona of a separable $\mathrm{C}^{*}$-algebras can be presented as a tensor product of infinite-dimensional $\mathrm{C}^{*}$-algebras ([83]).
11.0.2. The role of MA, $\mathrm{OCA^{\#}}$, and Biba’s trick